2.5. Determine the Cash Balance The beginning-of-the-period cash balance for the Travis Company L04 was a $12,000 debit. Cash sales for the month were $6,000 and sales on account were $8,000. The company paid $2,500 cash for current-period purchases and also paid $3,000 cash for amounts due from last month. What is the ending debit or credit balance in the Cash account?
Recording Transactions with the Accounting Equation During the year, the Decker Company expe rienced the following accounting transactions 1. Issued common stock in the amount of $100,000 2. Paid a $30,000 cash dividend 3. Borrowed $25,000 from a bank 4. Made a principal payment of $2,500 on an outstanding bank loan 5. Made an interest paynient of $1,200 on an outstanding bank loan LO2 Using the accounting equation, record each of the transactions in columnar format using the following template Assets = Liabilities CashNotes Payable +Common Stock Retained Earnings
Answer #1

Travis Company

Ending balance in the Cash account: $12500 debit

Cash
Beg. Bal. 12000 2500 Cash purchases
Cash sales 6000 3000 Payables
End. Bal. 12500

Decker Company

Assets = Liabilities + Stockholders' Equity
Cash = Notes Payable + Common Stock + Retained Earnings
1 100000 = + 100000 +
2 -30000 = + + -30000
3 25000 = 25000 + +
4 -2500 = -2500 + +
5 -1200 = + + -1200
Total 91300 = 22500 + 100000 + -31200